OpenStudy (anonymous):

The sum of two numbers is 15. The difference of the first number and three will equal the second number. Write and solve the system of equations using the substitution method.

OpenStudy (anonymous):

let the nos be x & y x+y=15 x-3=y x+x-3=15 2x=15+3 2x=18 x=18/2 or x=9 y=9-3=6 nos are 9,6